OpenStudy (anonymous):

Write an equation of the line, in point-slope , that passes through the two given points. points: (-2, 15), (9, -18)

OpenStudy (tkhunny):

Can you calculate the slope between the two points?

OpenStudy (noseboy908):

firstly, you need to find your slope.

OpenStudy (anonymous):

11/-33= -0.333 I think I did it wrong..

OpenStudy (tkhunny):

"I think I did it wrong." You DO know this makes no sense, right? Why not do it right, instead? \(Slope = \dfrac{Change\;in\;Y}{Change\;in\;X} = \dfrac{-18-15}{9-(-2)} = \dfrac{-33}{11} = -3\) You were upside down, that's all.